Communication Performance of Irregular Applications and the Implications for Graph Partitioners (extended abstract)

David R. O’Hallaron and Jonathan Richard Shewchuk

School of Computer Science Carnegie Mellon University Pittsburgh, PA 15213

Abstract

abstract goes here.

1 Introduction

2 Conclusions

A Error bound

This research was sponsored in part by the Advanced Research Projects Agency/CSTO monitored by SPAWAR under contract N00039-93-C-0152, in part by the National Science Foundation under Grant CMS-9318163, and in part by a grant from the Intel Corporation. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the of�cial policies, either express or implied, of the U.S. government.